MAA IRICA ASSCIATON Y ANERNG Logged in as aw webwork/math243spring-schmidt/ week 4c-_ch12 probablity/ Week 4c - Ch12 Probability: Problem 9 Previous Problen List Next (1 port) Suppose that a single die with 6 sides (numbered 1.2.3 rolls equals 7? Answer a 句isroled twice what istho probability that the sum of the two Preview My AnswersSubmit Answers You have attempted this problem O times You have unlimited attempts remaining Email instructor Page generabed at 0422018 an 112om POT WeBWorK O 1096-2016 themtdww.version: 2.12 pa.version 2.12 The eBwork Project
Answer #1

Solution :

Total possible ways to rolled a die twice is 6^2 = 36

The possible outcomes to the sum of the two rolls equals 7 is 6
{(1,6),(2,5),(3,4),(4,3),(5,2),(6,1)}

The probability that the sum of the two rolls equals to 7 is = possible outcomes/total possible ways
= 6/36 = 1/6
= 0.1667